OpenStudy (anonymous):

10. Fill in the missing reasons for the following two-column proof: Given: 5x + 1 = 21 Prove: x = 4 1) 5x + 1 = 21 1) Given 2) 5x = 20 2) _________________________ 3) x = 4 3) _________________________

OpenStudy (anonymous):

5x=21-1 5x=20 x=20/5 x=4

OpenStudy (anonymous):

thats understood but i need to write a two column proof

OpenStudy (wolf1728):

1) 5x + 1 = 21 1) Given 2) 5x = 20 2) Subtracting 1 from both sides 3) x = 4 3) Dividing both sides by 5

OpenStudy (anonymous):

you have to use one of the properties of equality to fill in a two column proof.

OpenStudy (wolf1728):

Something like 2) Subtracting the same number from both sides of an equation keeps both sides equal. ?

OpenStudy (wolf1728):

I found this link: http://hotmath.com/hotmath_help/topics/properties-of-equality.html So for step 2 Subtraction Property For all real numbers x, y, and z, if x = y, then x – z = y – z. This property allows you to balance and solve equations involving real numbers.

OpenStudy (wolf1728):

For step 3 Division Property For all real numbers x, y, and z, if x = y, and z ≠ 0, then x/z = y/z. This property allows you to balance and solve equations involving real numbers.

OpenStudy (anonymous):

okay then 3 would be...?

OpenStudy (wolf1728):

See my last posting

OpenStudy (anonymous):

im lost.

OpenStudy (anonymous):

because wouldnt that be for step 2

OpenStudy (wolf1728):

I don't know why you are lost. Step 2 pertains to subtraction and Step 3 pertains to division.

OpenStudy (wolf1728):

I can fill in the two column proof for all 3 steps if you want

OpenStudy (anonymous):

2 actually isnt substraction its division of 20/5 and 3 seems to be the solution by some property

OpenStudy (wolf1728):

1) 5x + 1 = 21 1) Given 2) 5x = 20 2) Subtraction Property For all real numbers x, y, and z, if x = y, then x – z = y – z. 3) x = 4 3) Division Property For all real numbers x, y, and z, if x = y, and z ≠ 0, then x/z = y/z. In step 2 we are subtracting 1 from both sides. In step 3 we are dividing both sides by 5

OpenStudy (anonymous):

but there was no subtraction done in 2. it was division...

OpenStudy (wolf1728):

Maybe it is better to add one more line to Step 2 1) 5x + 1 = 21 1) Given Then we subtract 1 from both sides 2) 5x = 20 2) Subtraction Property For all real numbers x, y, and z, if x = y, then x – z = y – z. We could add one more line to step 3 also

OpenStudy (anonymous):

but it would then be incorrect due to the fact there's only 3 lines lool

OpenStudy (wolf1728):

Okay time to add another instruction 1) 5x + 1 = 21 1) Given Then we subtract 1 from both sides 5x + 1 -1 = 21 -1 which gives us 2) 5x = 20 2) Subtraction Property For all real numbers x, y, and z, if x = y, then x – z = y – z.

OpenStudy (anonymous):

ohh okay

OpenStudy (anonymous):

i still dont get it though.

OpenStudy (wolf1728):

Well I have explained it about as much as I have. Do you understand about 2 column proofs?

OpenStudy (anonymous):

ughh i give up you get a medal for trying so hard!

OpenStudy (wolf1728):

Well thanks for the medal. I could explain even more but maybe I should quit. If you do find someone else to help you by all means direct them to this page.